a cook on a cruise ship is using a recipe that calls for 200 servings. the recipe needa 38 pounds of flour and and 110 cups of milk. The cook wnts to serve 500 people. How many pounds of flour and milk will he need?
Answer by rfer(16322)

200/38=500/x
200x=19000
x=19000/200
x=95 lbs flower
-----------------
200/110/500/x
200x=55000
x=55000/200
x=275 cups milk